OpenStudy (anonymous):

2-x(9+x^2) for x= -2

OpenStudy (nali):

just subtitute -2 in every place that has x and just solve

OpenStudy (anonymous):

Sub in x=-2. 2-x(9+x^2) =2-(-2)[9+(-2)^2]

OpenStudy (anonymous):

u can solve the expression first and then substitute -2 for all x's

OpenStudy (anonymous):

2-9(-2)-(-2)^3 =28